OpenStudy (anonymous):

given tan theta=7/(*radical*15) and sec theta > 0, find sin and cos theta

OpenStudy (anonymous):

this same as other one?

OpenStudy (anonymous):

no different

OpenStudy (anonymous):

draw a right triangle. label one angle theta, say. then put "opposite side" 7, "adjacent side" radical 15 solve for hypotenuse via pythagoras. you get \[h^2=7^2+\sqrt{15}^2=49+15=64\]so h = 8

OpenStudy (anonymous):

now you know all the ratios, so you are essentially done. for example \[\sin(\theta)=\frac{7}{15}\]